eviant:

Solve the equation using the Zero-Product Property. (x – 8)(5x – 6) = 0

eviant:

@dude

dude:

\((x – 8)(5x – 6) = 0\) Set the individual factors equal to 0 \(x-8=0\) and \(5x-6=0\) Solve for x on both

eviant:

@dude how do I solve for 5x-6?

dude:

Do the opposite of -6 on both sides Then do the opposite of multiplying 5 on both sides